# Asymptotic condition numbers for linear ODEs

### v1

@@ 3.4. Properties of the matrices $Q_{jl}(t)$
In our study of asymptotic forms and asymptotic condition numbers, the following properties of the matrices $Q_{jl}(t)$ in (42) are fundamental. We collect these properties in four propositions.
The first proposition relates the indices $j(u)$ and $L_{j}(u)$ to the condition $Q_{jl}(t)u=0$.
@@
    $Q_{jl}(t)u=0$ for $j\in\{1,\ldots,j(u)-1\}$ and $l\in\{0,\ldots,L_{j}\}$;
  • 3)
    $Q_{jl}(t)u=0$ for $j\in\{j(u),\ldots,q\}$ with $\Lambda_{j}(u)\neq\emptyset$ and $l\in\{L_{j}(u)+1,\ldots,L_{j}\}$.
In 2) and 3), we assume $u\neq 0$ in order to have $j(u)$ defined.
@@
For $j\in\{1,\ldots,q\}$ and $l\in\{0,\ldots,L_{j}\}$, we have
$Q_{jl}(t)u=\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P_{il}u$
by recalling (42).
Proof of point 1). Consider $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)=\emptyset$ and $l\in\{0,\ldots,L_{j}\}$. $\Lambda_{j}(u)=\emptyset$ implies, for any $\lambda_{i}\in\Lambda_{j}$, $\alpha^{(i)}(u)=0$ and then, by Propositions 12 and 13, for any $\lambda_{i}\in\Lambda_{j}$ we have $P_{ik}u=0$, $k\in\{0,\ldots,m_{i}-1\}$. In particular, for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$, we have $P_{il}u=0$.
Point 2) follows by point 1).
Proof of point 3). Consider $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)\neq\emptyset$ and $l\in\{L_{j}(u)+1,\ldots,L_{j}\}$. For any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)=0$, we have $P_{il}u=0$, since $P_{ik}u=0$, $k\in\{0,\ldots,m_{i}-1\}$. For any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)\neq 0$, we also have $P_{il}u=0$, since $P_{ik}u=0$, $k\in\{l_{i}(u)+1,\ldots,m_{i}-1\}$ by the definition of $l_{i}(u)$, and $l\geq L_{j}(u)+1\geq l_{i}(u)+1$. ∎
@@
Since (see (42))
$\left\|Q_{jl}(t)\right\|\leq\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\left\|P_{il}\right\|\text{\ \ and\ \ }\left\|Q_{jl}(t)u\right\|\leq\sum\limits_{\begin{subarray}{c}\lambda_{i}\in\Lambda_{j}\\ m_{i}\geq l+1\end{subarray}}\left\|P_{il}u\right\|$
hold, the points 1) and 2) follow.
Proposition 16 and Lemma 17, as applied to the linear combination
@@
where the second equality holds since for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)=0$ we have $P_{il}u=0$; and the third equality holds since for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$, $\alpha^{(i)}(u)\neq 0$ and $l_{i}(u)<l$ we have $P_{il}u=0$ by the definition of $l_{i}(u)$. Now, Proposition 15 and Lemma 17 imply 4). ∎
The third proposition says how the matrices $Q_{jl}(t)$ are transformed when we replace the matrix $A$ by $-A$. Observe that the matrix $-A$ has opposite eigenvalues, i.e. eigenvalues with opposite imaginary and real parts, with respect to the matrix $A$ and the dimensions of blocks and mini-blocks in the JCF of $-A$ are the same as in the JCF of $A$ (see Proposition 2 with $z=-1$). Therefore, we have the same number $q$ of different real parts for the eigenvalues of $-A$ and $A$. Moreover, the set $\Lambda_{j}(-A)$ and the numbers $r_{j}(-A)$ and $L_{j}(-A)$, $j\in\{1,\ldots,q\}$, corresponding to $-A$ are
@@
where $\Lambda_{q+1-j}$, $r_{q+1-j}$ and $L_{q+1-j}$ correspond to $A$. The indeces $j$ and $l$ for the matrices $Q_{jl}(t,-A)$ corresponding to $-A$ range over $j\in\{1,\ldots,q\}$ and $l\in\{0,\ldots,L_{q+1-j}\}$, respectively.
@@
Now, use Proposition 9 with $z=-1$. ∎
In the case of a real matrix $A$, the fourth proposition explains how to rewrite the expression (42) that defines $Q_{jl}(t)$ in terms of the real eigenvalues and complex conjugate pairs of eigenvalues of $A$. For a matrix $Z$, we denote by $\mathrm{Re}(Z)$ the matrix given by the real parts of the elements of $Z$.
@@
For $\lambda_{i}\in\Lambda_{j}$ such that $\lambda_{i}$ is real, we have $\omega_{i}=0$. For a complex conjugate pairs $\lambda_{i_{1}},\lambda_{i_{2}}\in\Lambda_{j}$ with $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ and $\omega_{i_{1}}>0$, by Proposition 10 we have
@@
has zero or one term $P_{il}$, which is real (see Proposition 10). Moreover, in the other sum, each term
@@
3.5. The matrices $Q^{e}_{j}(t)$
The contents of this section are used for defining in Subsection 4.6 the global asymptotic condition number of the problem (2).
@@ 3.5. The matrices $Q^{e}_{j}(t)$
the matrices of the (right) eigenvectors and left eigenvectors, respectively, corresponding to $\lambda_{i}$ (see Section 2). The superscript or subscript $e$ in the notation denotes “eigen”.
Let $U^{e}_{j}$ be the sum of the eigenspaces corresponding to eigenvalues in $\Lambda_{j}$. Observe that the linear operator $Q^{e}_{j}(t)$ maps $\mathbb{C}^{n}$ in the subspace $U^{e}_{j}$. Therefore, $Q^{e}_{j}(t)|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$. The next proposition states that the restriction $Q^{e}_{j}(t)|_{U^{e}_{j}}$ is invertible.
@@
$Q^{e}_{j}(-t)Q^{e}_{j}(t)$ $=$ $\left(\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{-\sqrt{-1}\ \omega_{i}t}P^{e}_{i}\right)\left(\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{i}t}P^{e}_{i}\right)$
$=$ $\sum\limits_{\lambda_{i}\in\Lambda_{j}}P^{e}_{i}$
and then
@@
  • 2.
    In the case where $\Lambda_{j}$ consists of a real eigenvalue $\lambda_{i}$, we have
    $Q^{e}_{j}(t)|_{U^{e}_{j}}=P_{i}^{e}|_{U^{e}_{j}}=I_{U^{e}_{j}}.$
Next proposition considers the linear operators $P_{i}^{e}|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$, $\lambda_{i}\in\Lambda_{j}$.
@@
$0=\sum\limits_{i=1}^{p}c_{i}P_{i}^{e}|_{U^{e}_{j}}.$
Consider $u\in U^{e}_{j}$ such that $\alpha_{ij^{\prime}1}(u)\neq 0$ for $\lambda_{i}\in\Lambda_{j}$ and $j^{\prime}\in\{1,\ldots,d_{i}\}$. Since $u\in U^{e}_{j}$, we have $\alpha_{ij^{\prime}k}(u)=0$ for $\lambda_{i}\in\Lambda_{j}$, $j^{\prime}\in\{1,\ldots,d_{i}\}$ and $k\in\{2,\ldots,m_{ij^{\prime}}\}$. Thus $l_{i}(u)=0$ for $\lambda_{i}\in\Lambda_{j}$, by Proposition 14. Since (recall Subsection 2.3.1)
$P_{i}^{e}u=P_{i0}u,\ \lambda_{i}\in\Lambda_{j},$
@@
where we set $c_{i0}=c_{i}$, $i\in\{1,\ldots,p\}$. By using Proposition 15, we obtain $c_{i}=c_{i0}=0$, $\lambda_{i}\in\Lambda_{j}$. ∎
@@
As a consequence of the previous proposition and Lemma 17, we obtain

### v2

@@ Appendix B Properties of the matrices $Q_{jl}(t)$
In our study of asymptotic forms and asymptotic condition numbers, the following properties of the matrices $Q_{jl}(t)$ defined in (12) are fundamental. We collect these properties in four propositions.
The first proposition relates $\Lambda_{j}(u)$ and $L_{j}(u)$, defined in Subsection 2.4, to the condition $Q_{jl}(t)u=0$.
@@
    $Q_{jl}(t)u=0$ for $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)\neq\emptyset$ and $l\in\{L_{j}(u)+1,\ldots,L_{j}\}$.
@@
Proof of point 1). Consider $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)=\emptyset$ and $l\in\{0,\ldots,L_{j}\}$. $\Lambda_{j}(u)=\emptyset$ implies, for any $\lambda_{i}\in\Lambda_{j}$, $\alpha^{(i)}(u)=0$ and then, for any $\lambda_{i}\in\Lambda_{j}$ we have $P_{ik}u=0$, $k\in\{0,\ldots,m_{i}-1\}$, by Propositions 33 and 34 in Appendix A. In particular, we have $P_{il}u=0$ in (12), for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$.
Proof of point 2). Consider $j\in\{1,\ldots,q\}$ with $\Lambda_{j}(u)\neq\emptyset$ and $l\in\{L_{j}(u)+1,\ldots,L_{j}\}$. In (12), for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)=0$, we have $P_{il}u=0$, since $P_{ik}u=0$, $k\in\{0,\ldots,m_{i}-1\}$. Moreover, for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)\neq 0$, we also have $P_{il}u=0$, since $P_{ik}u=0$, $k\in\{l_{i}(u)+1,\ldots,m_{i}-1\}$ by the definition of $l_{i}(u)$ given in Appendix A, and $l\geq L_{j}(u)+1\geq l_{i}(u)+1$. ∎
@@
The points 1) and 2) are trivial: by (12), we have
[…]
Proposition 38 and Lemma 39 in Appendix A, as applied to the linear combination
@@
where the second equality holds since, for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$ and $\alpha^{(i)}(u)=0$, we have $P_{il}u=0$ by Propositions 33 and 34 in Appendix A; and the third equality holds since, for any $\lambda_{i}\in\Lambda_{j}$ with $m_{i}\geq l+1$, $\alpha^{(i)}(u)\neq 0$ and $l_{i}(u)<l$, we have $P_{il}u=0$ by the definition of $l_{i}(u)$ in Appendix A. Now, Proposition 37 and Lemma 39 in Appendix A imply 4). ∎
The third proposition says how the matrices $Q_{jl}(t)$ are transformed when we replace the matrix $A$ by A-A. Observe that the matrix A-A has opposite eigenvalues, i.e., eigenvalues with opposite imaginary and real parts, with respect to the matrix $A$ and the dimensions of blocks and mini-blocks in the JCF of A-A are the same as in the JCF of $A$: see Proposition 21 in Appendix A with z=1z=-1. Therefore, we have the same number $q$ of different real parts for the eigenvalues of A-A and $A$. Moreover, the set $\Lambda_{j}(-A)$ and the numbers $r_{j}(-A)$ and $L_{j}(-A)$, $j\in\{1,\ldots,q\}$, corresponding to A-A are
@@
where $\Lambda_{q+1-j}$, $r_{q+1-j}$ and $L_{q+1-j}$ correspond to $A$. The indices $j$ and $l$ for the matrices $Q_{jl}(t,-A)$ corresponding to A-A range over $j\in\{1,\ldots,q\}$ and $l\in\{0,\ldots,L_{q+1-j}\}$, respectively.
@@
Now, use Proposition 29 in Appendix A with z=1z=-1. ∎
In the case of a real matrix $A$, the fourth proposition explains how to rewrite the expression (12) that defines $Q_{jl}(t)$ in terms of the real eigenvalues and complex conjugate pairs of eigenvalues of $A$. For a matrix $Z$, we denote by $\mathrm{Re}(Z)$ and $\mathrm{Im}(Z)$ the matrices given by the real parts and imaginary parts, respectively, of the elements of $Z$.
@@
For $\lambda_{i}\in\Lambda_{j}$ such that $\lambda_{i}$ is real, we have $\omega_{i}=0$. For a complex conjugate pair $\lambda_{i_{1}},\lambda_{i_{2}}\in\Lambda_{j}$ with $\lambda_{i_{2}}=\overline{\lambda_{i_{1}}}$ and $\omega_{i_{1}}>0$, by Proposition 30 in Appendix A we have
@@
has zero or one term $P_{il}$, which is real: see Proposition 30 in Appendix A. Moreover, in the other sum, each term
@@
Appendix C The matrices $Q^{e}_{j}(t)$
The contents of this section are used for defining in Subsections 4.2 and 4.3 the global asymptotic condition number $K_{\infty}(t)$ of the problem (2).
@@ Appendix C The matrices $Q^{e}_{j}(t)$
the matrices of the (right) eigenvectors and left eigenvectors, respectively, corresponding to $\lambda_{i}$ (see Appendix A). The superscript, or subscript, $e$ in the notation stands for “eigen”.
Let $U^{e}_{j}$ be the sum of the eigenspaces corresponding to eigenvalues in $\Lambda_{j}$. Observe that the linear operators $P_{i}^{e}$, $\lambda_{i}\in\Lambda_{j}$, and $Q^{e}_{j}(t)$ map $\mathbb{C}^{n}$ into the subspace $U^{e}_{j}$. Therefore, $P_{i}^{e}|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$ and $Q^{e}_{j}(t)|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$.
Remark 45.
When $\Lambda_{j}$ consists of a real eigenvalue $\lambda_{i}$, we have
$Q^{e}_{j}(t)|_{U^{e}_{j}}=P_{i}^{e}|_{U^{e}_{j}}=I_{U^{e}_{j}}.$
The next proposition establishes that $P_{i}^{e}$ and $Q_{j}^{e}(t)$ coincide with $P_{i0}$ and $Q_{j0}(t)$ when restricted to $U_{j}^{e}$. $P_{i0}$ is defined in (46) of Appendix A (see also Subsection A.3.1 in Appendix A) and $Q_{j0}$ is defined in (12).
Proposition 46.
Let $j\in\{1,\ldots,q\}$. We have $P_{i}^{e}|_{U^{e}_{j}}=P_{i0}|_{U^{e}_{j}}$, $\lambda_{i}\in\Lambda_{j}$, and $Q^{e}_{j}(t)|_{U^{e}_{j}}=Q_{j0}(t)|_{U^{e}_{j}}$.
Proof.
For $\lambda_{i}\in\Lambda_{j}$, $P_{i}^{e}|_{U^{e}_{j}}=P_{i0}|_{U^{e}_{j}}$ follows by Subsection A.3.1 in Appendix A and the fact that, for $u\in U^{e}_{j}$, we have $\alpha_{ij^{\prime}k}(u)=0$, for $\lambda_{i}\in\Lambda_{j}$, $j^{\prime}\in\{1,\ldots,d_{i}\}$ and $k\in\{2,\ldots,m_{ij^{\prime}}\}$. $Q^{e}_{j}(t)|_{U^{e}_{j}}=Q_{j0}(t)|_{U^{e}_{j}}$ follows by $P_{i}^{e}|_{U^{e}_{j}}=P_{i0}|_{U^{e}_{j}}$, $\lambda_{i}\in\Lambda_{j}$. ∎
The next proposition states that the restriction $Q^{e}_{j}(t)|_{U^{e}_{j}}$ is invertible.
@@
$Q^{e}_{j}(-t)Q^{e}_{j}(t)$ $=$ $\left(\sum\limits_{\lambda_{i}\in\Lambda_{j}}\mathrm{e}^{-\sqrt{-1}\ \omega_{i}t}P^{e}_{i}\right)\left(\sum\limits_{\lambda_{k}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ \omega_{k}t}P^{e}_{k}\right)$
$=$ $\sum\limits_{\lambda_{i},\lambda_{k}\in\Lambda_{j}}\mathrm{e}^{\sqrt{-1}\ (-\omega_{i}+\omega_{k})t}P^{e}_{i}P^{e}_{k}=\sum\limits_{\lambda_{i}\in\Lambda_{j}}P^{e}_{i},$
since $P^{e}_{i}P^{e}_{k}=0$ for $\lambda_{i}\neq\lambda_{k}$ and $P^{e}_{i}P^{e}_{k}=P^{e}_{i}$ for $\lambda_{i}=\lambda_{k}$. Thus,
@@
Next proposition considers the linear independence of the linear operators $P_{i}^{e}|_{U^{e}_{j}}:U^{e}_{j}\rightarrow U^{e}_{j}$, $\lambda_{i}\in\Lambda_{j}$.
@@
$0=\sum\limits_{\lambda_{i}\in\Lambda_{j}}c_{i}P_{i}^{e}|_{U^{e}_{j}}.$
Consider $u\in U^{e}_{j}$ such that $\alpha_{ij^{\prime}1}(u)\neq 0$ for $\lambda_{i}\in\Lambda_{j}$ and $j^{\prime}\in\{1,\ldots,d_{i}\}$. Since $u\in U^{e}_{j}$, we have $\alpha_{ij^{\prime}k}(u)=0$ for $\lambda_{i}\in\Lambda_{j}$, $j^{\prime}\in\{1,\ldots,d_{i}\}$ and $k\in\{2,\ldots,m_{ij^{\prime}}\}$. Thus $l_{i}(u)=0$ for $\lambda_{i}\in\Lambda_{j}$, by Proposition 36 in Appendix A.
@@
where the second equality follows by Proposition 46 in this appendix, and we set $c_{i0}=c_{i}$, $\lambda_{i}\in\Lambda_{j}$. By using Proposition 37 in Appendix A, we obtain $c_{i}=c_{i0}=0$, $\lambda_{i}\in\Lambda_{j}$. ∎
@@
As a consequence of the previous Proposition 49 and Lemma 39 in Appendix A, we obtain